OpenStudy (anonymous):

The side of an equilateral triangle is 3 inches shorter than the side of a square. The perimeter of the square is 37 inches more than the perimeter of the triangle. Find the length of a side of the square.

OpenStudy (lgbasallote):

let x = triangle side y = square side x = y -3 4y = 37 + 3x 4y = 37 + 3(y - 3) 4y = 37 + 3y - 9 y = 28